OpenStudy (anonymous):

solution of system -10 + 6x + 7y = 0 and -x + 2 = y. (solve with matrices). i got x = 4 and y = -2, is that correct?

OpenStudy (jim766):

nope, that is not what I got, how did you set up your matrices?

OpenStudy (anonymous):

@Jim766 my friend helped me do it so i have no idea. he just gave me the answer he got. i tried to find how to set up the matrices and it totally lost me. how did you find it?

OpenStudy (jim766):

to set up the matrices you need the x's and y's on one side and the plain numbers on the other -10 + 6x + 7y = 0 +10 +10 6x + 7y = 10 do you see what I did?

OpenStudy (anonymous):

okay that makes sense, but how do you do it with -x + 2 = y, because the variables are on both sides of the equatio

OpenStudy (jim766):

-x +2 = y -2 -2 -x = y -2 -y -y -x - y = -2

OpenStudy (jim766):

I gotta tell you that I screwed up when I did it, your answer was correct, do you still want to see how to do it?

OpenStudy (anonymous):

i think i get it now. my matrix will be [6 7; -1 1] x [x; y] = [10; -2], right?

OpenStudy (anonymous):

i had no idea how to do the matrices in the equation thing lol

OpenStudy (jim766):

right are you good from there?

OpenStudy (anonymous):

i'm still not entirely sure how to solve for x and y with a matrix, even though the answer my friend found was right.

OpenStudy (jim766):

are you using a graphing calculator?

OpenStudy (anonymous):

i could if i need to. would it be easier than solving it on paper?

OpenStudy (jim766):

yeah, way easier.... what kind do you have?

OpenStudy (anonymous):

ti84 plus

OpenStudy (jim766):

do you know how to enter in a matrix?

OpenStudy (jim766):

get your calculator

OpenStudy (anonymous):

i'm trying to figure out why it wont let me put a negative number in the matrix

OpenStudy (anonymous):

got it.

OpenStudy (jim766):

it will .... start from the beginning here are the equations 6x + 7y = 10 -1x -1y = -2 here is my one matrix 6 7 10 -1 -1 -2 see how I got that?

OpenStudy (anonymous):

yep

OpenStudy (jim766):

push 2nd matrix and slide over to edit make the matrix 2X3

OpenStudy (anonymous):

i have all the values in, now what?

OpenStudy (jim766):

now save it by pressing 2nd quit

OpenStudy (jim766):

now press 2nd matrix and slide to the middle (math)

OpenStudy (jim766):

go up to B: rref( press enter press2nd matrix again and enter

OpenStudy (jim766):

should look like this rref( [A] press enter

OpenStudy (jim766):

1 0 4 0 1 -2

OpenStudy (anonymous):

the last 2 values are my answers. whaaaaaaaat it's like witchcraft

OpenStudy (jim766):

right it translates to 1x + 0y = 4 x = 4 and 0x + 1y = -2 y = -2 the point where they intersect is 4, 2

OpenStudy (jim766):

it is pretty magical...

OpenStudy (anonymous):

that's awesome. thank you for your help!

OpenStudy (jim766):

np yw